Bed & Bath, a retailing company, has two departments—Hardware and Linens. The company’s most recent monthly contribution format income statement follows:

Department
Total Hardware Linens
Sales $ 4,050,000 $ 3,040,000 $ 1,010,000
Variable expenses 1,330,000 923,000 407,000
Contribution margin 2,720,000 2,117,000 603,000
Fixed expenses 2,130,000 1,310,000 820,000
Net operating income (loss) $ 590,000 $ 807,000 $ (217,000 )

A study indicates that $372,000 of the fixed expenses being charged to Linens are sunk costs or allocated costs that will continue even if the Linens Department is dropped. In addition, the elimination of the Linens Department will result in a 18% decrease in the sales of the Hardware Department.

Required:

What is the financial advantage (disadvantage) of discontinuing the Linens Department?
Answer #1

Contribution margin for Hardware would be=2,117,000*(1-Growth Rate)

=2,117,000*(1-0.18)=$1735940

Less:Fixed cost for Hardware=1,310,000

Net operating income for Hardware=425940

Less:Unavoidable fixed cost for Linens=372,000

Total net operating income=$53940

Hence financial disadvantage=590,000-53940

=$536060